OpenStudy (anonymous):

Given: x + y = 6. If x € {-15, -6, -1}, then which of the following sets of ordered pairs are solutions? {(-15, -21), (-6, -12), (-1, -7)} {(-15, 21), (-6, 12), (-1, 7)} {(-15, 21), (-6, 12), (-1, -7)}

OpenStudy (anonymous):

x+y=8 y=6-x {y}=6-{x} {y}=6-{-15,-6,-1} {y}={6-(-15),6-(-6),6-(-1)} {y}={21,12,7} B